Find the vertex and the axis of symmetry of the parabola:
-4x^2+2x+11=0

Answer by Alan3354(69443)

Find the vertex and the axis of symmetry of the parabola:
f(x) = -4x^2+2x+11=0
---------
The AOS is x = -b/2a
-----
Vertex is (x,f(x)), x = -b/2a